OpenStudy (anonymous):

describe the surface whose equation is given as y=x^2

OpenStudy (anonymous):

(x+5)^2 + (y+2)^2 + (z+1)^2 = 19+25+4+1 = 49 sphere C(-5,-2,-1) R = 7